# LAMD: Context-driven Android Malware Detection and Classification with LLMs

### v1

@@ 1 Introduction
  • We evaluate LAMD on a collected dataset [^1] The dataset is open-source for further research on LLM-based malware tasks: https://doi.org/10.5281/zenodo.14884736 that reflects the real-world setting, demonstrating its effectiveness in detecting and explaining Android malware, outperforming conventional detectors.
@@ 2.1 Learning-based Android Malware Detection
Learning-based Android malware detectors leverage machine learning or deep learning models to automatically learn patterns from features extracted by static or dynamic analysis, enhancing scalability and adaptability [10]. Due to the high cost of dynamic analysis, most models rely on static feature extraction through reverse engineering [11, 12]. Despite advancements, these models struggle with concept drift, where evolving malware variants degrade performance in real-world deployments [1, 2]. Current works try to mitigate it, some of them focus on exploring robust features [13, 14, 15, 16] and others leverage continual learning [17, 3, 18] or active learning [19, 20] to let models adapt to new distribution. However, these methods either target to specific feature space [13] or introduce high retraining overhead and the risk of label poisoning [21, 22]. An additional challenge is explainability, which is critical for security analysis. Existing methods primarily use feature attribution techniques, providing importance scores without generating human-readable behavioral analysis [23, 24]. To address these limitations, researchers are increasingly exploring LLM-based approaches, which leverage extensive external knowledge and reasoning capabilities to improve malware detection and analysis [25, 26].
@@ 3.2.2 Backward Program Slicing
Extracting suspicious APIs is useful, but analyzing them in isolation often obscures their context. For instance, while sendTextMessage() is legitimate in messaging apps, malware may exploit it for premium-rate SMS. To capture intent, we extract functions invoking suspicious APIs and refine their control flow graphs (CFGs), $G=<N,E>$, where nodes $n\in N$ represent basic blocks or instructions, and edges $(e_{1},e_{2})\in E$ define control flow. Since CFGs can be large and noisy, we apply backward slicing [40] to isolate instructions affecting the API invocation. A slice $S$ is defined by a slicing criterion $C=<s,V>$ where $s$ is the statement invoking $a_{i}$ and $V$ includes all parameters. We classify relevant variables as: (1) Direct relevant variables: Variables’ values can affect variable $v\in V$ of $a_{i}$ (2) Indirect relevant variables: Variables in branch statements whose value affects invocation of $a_{i}$. The backward slicing is to select the set of instructions in $\mathcal{P}$ that directly or indirectly affect the execution or parameters of $a_{i}$. The backward slicing algorithm consists of two steps to ensure completeness in complex branch structures:
@@
After slicing, we generate sliced CFGs for each sensitive API, preserving essential control flow and statements. Notably, If undeclared variables remain in a sliced function, inter-procedural backward slicing is recursively applied to its callers until all variables are resolved. The details of slicing algorithm are shown in Appendix E.
@@ 4 Evaluation
This section presents a comprehensive evaluation of LAMD, assessing its Android malware detection performance in real-world scenarios and the quality of its generated explanations through a series of experiments. Experiment setup details are shown in Appendix A.
4.1 Dataset Construction
Based on these principles, we select Android APKs from Androzoo [44] according to their discovery time, which is determined by their submission to VirusTotal [45] (as release timestamps can be unreliable [^2] Due to modifiable or randomly generated release times, some timestamps (dex date) are inaccurate.). The dataset spans from 2014 to 2023, comprising 13,794 samples. The remaining samples are evenly split into three test sets (about 3,015 each) to represent incremental distribution drift. Details of the dataset, including its source, drift severity, and training size considerations, are provided in Appendix B.
4.2 Metric
@@
{tblr}
cells = c, colsep=2.5pt, cell11 = r=2, cell12 = c=3, cell15 = c=3, cell18 = c=3, cell41 = valign=m, vline2-3,6 = 1, vline5,8 = 2, vline2,5,8 = 1-8, hline1,3-9 = -, hline2 = 2-10, Model & Test 1 Test 2 Test 3
@@
4.3 Baseline
In Android malware detection, an APK serves as input, containing the codebase (e.g., .dex files) and configuration files (e.g., AndroidManifest.xml), which provide behavioral insights like API calls and permissions, represented in vector or graph formats. We evaluate LAMD against Drebin [11], DeepDrebin [12], and Malscan [52] which are representative learning-based methods on these feature formats, with details in Appendix D. To assess component impact, LAMD-R removes tier-wise reasoning to test structural and semantic analysis, while LAMD-F retains reasoning but omits factual consistency verification to evaluate hallucination control.
4.4 Evaluation Results
@@ 4.4.1 Malware Detection Performance
Table I compares LAMD’s detection performance with baselines. LAMD improves F1-scores by 23.12% and reduces FNR by 71.59% on average, enhancing detection reliability. While the FPR shows a slight increase, it is less indicative of true performance due to class imbalance, where learning-based methods often misclassify malware as benign because of the dominance of benign sampls. The performance drop in LAMD-R highlights the necessity of hierarchical code summarization, and the slight decline in LAMD-F underscores the role of hallucination mitigation. By refining input code and extracting key information, LAMD minimizes hallucination risks, ensuring more reliable malware detection.
@@ 4.4.2 Effectiveness of explanations
To assess analysis quality, we validate 100 correctly detected malware samples. Results show that 81 out of 100 samples are correctly classified into their respective categories. Table II summarizes the sample distribution and classification accuracy across categories. Due to their less distinct malicious patterns, Adware and Riskware pose greater challenges to accurate analysis compared to other categories, contributing to their higher misclassification rates.
TABLE II: The quality evaluation of generated detection analysis. The “Correct” column indicates the number of categories correctly classified by the LLM.
@@
PUA
umpay, scamapp, apptrack
13 9
@@ 5 Case Study
This section demonstrates how LAMD overcomes the limitations of current LLMs in Android malware detection. Since LLMs cannot process APKs directly, we use JADX [^3] https://github.com/skylot/jadx to decompile them, concatenating all pesudo source codes for input [47]. Besides GPT-4o-mini, we select Gemini 1.5 pro [6] as another comparison, which claims their longest context windows and for its malware detection capabilities.
@@ 6 Discussion
While LLMs improve Android malware detection, their general pre-training limits fine-grained behavior analysis. Future research should enhance security expertise through domain-specific fine-tuning or external knowledge integration. Meanwhile, learning-based methods remain essential for capturing overall insights in non-drift scenarios, highlighting effective strategies to combine these two paradigms and maximize their complementary strengths.
@@ 7 Conclusion
Appendix A Experiment Details
Appendix B Dataset Description
B.1 Dataset Source
We construct a dataset spanning 2014 to 2023666as APK labels generally stabilize after about one year in the wild [53], we choose samples before 2024 using samples from the Androzoo [44] platform, a comprehensive repository aggregating samples from sources such as Google Play, PlayDrone, VirusShare, and AppChina. Each sample’s timestamp corresponds to its submission date on VirusTotal [45], representing the time it was discovered. The dataset includes a decade of benign and malicious samples, labeled based on VirusTotal analysis from Androzoo. Samples flagged by more than four vendors are classified as malicious. To track malware family evolution, we used Euphony [54] to extract family labels. Table shows details about the used dataset.
TABLE III: Overview of the evaluation dataset, where M denotes malware and B denotes benign applications.
Time Interval Sample Size
The number of
Existing family
The number of
New family
Packed Malicious Benign M/(M+B)%
Test set 1 2020.05 - 2021.01 3015 21 24 18 284 2731 9.42
Test set 2 2021.01 - 2021.12 3015 28 32 30 298 2717 9.88
Test set 3 2021.12 - 2023.12 3016 34 36 40 302 2714 10.01
TABLE IV: Performance of learning-based methods within Drebin feature under different training/validation data volumes.
{tblr}
cells = c, cell11 = r=2, cell12 = r=2, cell13 = c=3, cell16 = c=3, cell19 = c=3, cell112 = c=3, cell31 = r=2, cell51 = r=2, vline2-3,6,9,12 = 1-7, vline3,6,9,12 = 3,5, vline3,6,9,12 = 4,6, hline1,3,5,7 = -, hline2 = 3-14, hline4,6 = 2-14, Sample Size & Model Validation Test 1 Test 2 Test 3
@@ B.1 Dataset Source
B.2 Dataset Drift and Training Size
@@
Intuitively, increased training samples should enhance detection performance in learning-based methods. While LAMD excels in zero-shot learning, one may argue that traditional methods are hindered by smaller training datasets compared to the vast pretraining data leveraged by LLMs. However, the impact of significant drift warrants reconsideration. Table IV highlights the performance of learning-based detectors (e.g., Drebin feature space) across varying dataset sizes, across the training dataset timeframe. The results indicate that additional training data does not improve detection of samples with significant drift, though it enhances performance on in-distribution samples.
@@
Appendix C Data Dependency used for Factual Consistency Verification
To check factual consistency, we leverage data dependencies that capture relationships between variables and APIs. We focus on five dependencies (see Table V) as they represent fundamental program relationships critical for malware reasoning and detection. Specifically:
@@ Appendix C Data Dependency used for Factual Consistency Verification
TABLE V: Data dependencies are used for factual consistency verification.
{tblr}
cell21 = r=3, cell51 = r=2, vline2-3 = 1-4,5-6, vline3 = 3-4,6, hline1-2,5,7 = -, hline3-4,6 = 2-3, Category & Type Description
Transitive A variable is propagated through a call chain to an API.
Derived A variable is derived from another through computation.
Appendix D Baseline Models
The details of baseline models are shown as follows:
  • Drebin: Drebin [11] detects malware using binary feature vectors derived from nine data types (e.g., hardware components, API calls, permissions) and classifies samples via a linear classifier.
  • DeepDrebin: DeepDrebin [12] extends Drebin by replacing the linear classifier with a three-layer deep neural network (DNN) while retaining the same feature space for feature extraction and classification.
  • Malscan: Malscan [52] employs a graph-based approach, extracting sensitive API calls from APKs and computing four centrality measures (degree, Katz, proximity, and harmonic wave centralities) as features. We use the optimal feature with a Random Forest classifier, which leads to the lowest overhead.
Appendix E Backward Slicing Algorithm

### v2

@@ 1 Introduction
  • We evaluate LAMD on a collected dataset [^1] The dataset is open-source for further research on LLM-based malware tasks: https://doi.org/10.5281/zenodo.14884736 that reflects the real-world setting. Our results show that LAMD outperforms conventional learning-based detectors under distribution shift, demonstrating its effectiveness in detecting and explaining fast evolving Android malware. Furthermore, we discuss the tension between learning-based models and pre-trained LLMs, highlighting promising directions for future research.
@@ 2.1 Learning-based Android Malware Detection
Learning-based Android malware detectors leverage machine learning or deep learning models to automatically learn patterns from features extracted by static or dynamic analysis, enhancing scalability and adaptability [10]. Due to the high cost of dynamic analysis, most models rely on static feature extraction through reverse engineering [11, 12]. Despite advancements, these models struggle with concept drift, where evolving malware variants degrade performance in real-world deployments [1, 2]. Current works try to mitigate it, some of them focus on exploring robust features [13, 14, 15, 16] and others leverage continual learning [17, 3, 18] or active learning [19, 20] to let models adapt to new distribution. However, these methods either target to specific feature space [13] or introduce high retraining overhead and the risk of label poisoning [21, 22]. An additional challenge is explainability, which is critical for security analysis. Existing methods primarily use feature attribution techniques, providing importance scores without generating human-readable behavioral analysis [23, 24]. To address these limitations, researchers are increasingly exploring LLM-based approaches, which leverage extensive external knowledge and reasoning capabilities [25, 26], while tasks related to malware detection and analysis still remain under-explored.
@@ 3.2.2 Backward Program Slicing
Extracting suspicious APIs is useful, but analyzing them in isolation often obscures their context. For instance, while sendTextMessage() is legitimate in messaging apps, malware may exploit it for premium-rate SMS. To capture intent, we extract functions invoking suspicious APIs and refine their control flow graphs (CFGs), $G=<N,E>$, where nodes $n\in N$ represent basic blocks or instructions, and edges $(e_{1},e_{2})\in E$ define control flow.
Recent work reveals that decoder-only Transformers, at the core of modern LLMs like GPT-4o-mini, are susceptible to over-squashing [40], where information from distant tokens is compressed and loses influence. The authors show how this may impair models in learning proper representations of similar long sequences, leading to representation clashes and wrong outcomes for downstream tasks. The study suggests that input-compressing strategy might preserve a high the signal-to-noise ratio, able to improve performance on downstream tasks. We follow this intuition in our context (the analysis of large complex programs), aiming at providing quality context to the models to improve on the downstream tasks. Since CFGs can be large and noisy, containing instructions that are irrelevant to suspicious API usages, directly inputting them to LLM risks conflating malicious and benign patterns, thereby obscuring sensitive behaviors. We therefore apply backward slicing [41] to isolate instructions affecting the API invocation and provide meaningful context to LLM. A slice $S$ is defined by a slicing criterion $C=<s,V>$ where $s$ is the statement invoking $a_{i}$ and $V$ includes all parameters. We classify relevant variables as: (1) Direct relevant variables: Variables’ values can affect variable $v\in V$ of $a_{i}$ (2) Indirect relevant variables: Variables in branch statements whose value affects invocation of $a_{i}$. The backward slicing is to select the set of instructions in $\mathcal{P}$ that directly or indirectly affect the execution or parameters of $a_{i}$. The backward slicing algorithm consists of two steps to ensure completeness in complex branch structures:
@@
After slicing, we generate sliced CFGs for each sensitive API, preserving essential control flow and statements. Notably, If undeclared variables remain in a sliced function, inter-procedural backward slicing is recursively applied to its callers until all variables are resolved. The details of slicing algorithm are shown in Appendix D.
@@ 4 Evaluation
Model Test 1 Test 2 Test 3
@@
This section presents a comprehensive evaluation of LAMD, assessing its Android malware detection performance in real-world scenarios and the quality of its generated explanations through a series of experiments.
4.1 Dataset Construction
Based on these principles, we select Android APKs from Androzoo [47] according to their discovery time, which is determined by their submission to VirusTotal [48] (as release timestamps can be unreliable [^2] Due to modifiable or randomly generated release times, some timestamps (dex date) are inaccurate.). The dataset includes a decade of benign and malicious samples, labeled based on VirusTotal analysis from Androzoo. Samples flagged by more than four vendors are classified as malicious. To track malware family evolution, we used Euphony [49] to extract family labels. The dataset spans from 2014 to 2023333as APK labels generally stabilize after about one year in the wild [50], we choose samples before 2024, comprising 13,794 samples. The remaining samples are evenly split into three test sets (about 3,015 each) to represent incremental distribution drift. Table VI shows details about the used dataset.
4.2 Metrics
4.3 Baselines
In Android malware detection, an APK serves as input, containing the codebase (e.g., .dex files) and configuration files (e.g., AndroidManifest.xml), which provide behavioral insights like API calls and permissions, represented in vector or graph formats. We evaluate LAMD against Drebin [11], DeepDrebin [12], and Malscan [45] which are representative learning-based methods on these feature formats, with details in Appendix A. To assess component impact, LAMD-R removes tier-wise reasoning to test structural and semantic analysis, while LAMD-F retains reasoning but omits factual consistency verification to evaluate hallucination control.
TABLE II: Out-of-distribution (OOD) detection performance (%) on post-cutoff samples for LAMD and aligned training-temporal scope for learning-based models.
Model F1 FPR FNR
Drebin [11] 83.72 0.42 22.02
DeepDrebin [12] 79.78 0.39 26.80
Malscan [45] 83.87 0.00 46.77
LAMD 85.71 1.89 14.29
4.4 Experiments
We conduct a series of experiments to thoroughly evaluate LAMD’s performance in malware detection and the quality of its generated explanation.
@@ 4.4.1 Malware Detection Performance
Table I compares LAMD’s detection performance with baselines on test datasets with the increasing level of drift. LAMD improves F1-scores by 23.12% and reduces FNR by 71.59% on average, enhancing detection reliability. While the FPR shows a slight increase, it is less indicative of true performance due to class imbalance, where learning-based methods often misclassify malware as benign because of the dominance of benign samples. The performance drop in LAMD-R highlights the necessity of hierarchical code summarization, and the slight decline in LAMD-F underscores the role of hallucination mitigation. By refining input code and extracting key information, LAMD minimizes hallucination risks, ensuring more reliable malware detection.
We performed an additional out-of-distribution (OOD) experiments designed to simulate a temporal natural drift by aligning GPT-4o-mini’s pre-training cutoff date (October 2023) with a dataset comprising samples of newer timestamps. As detailed in Table V, to make a fair comparison, LAMD is tested on samples from November and December 2023, while learning-based models are assessed from May-June 2020, the first two months following their training phase. Table II confirms the trend we have seen in Table I: LAMD outperforms baselines and state-of-the-art methods even under this fair settings. We also observe notable tension between the two paradigms: the breakdown of F1 score reports a higher FPR and a lower FNR for LAMD compared to learning-based approaches. We hypothesize that zero-shot learners such as LAMD have better generalizability benefit from their large-scale pre-trained dataset that may generalize better against similar threats (resulting in lower FNR). Conversely, they suffer from making more mistakes due to the lack of a smaller, task-dependent dataset (lower FPR for learning-based approaches). If true, this opens interesting research opportunities to study tensions between the two paradigms.
@@ 4.4.2 Effectiveness of explanations
To assess analysis quality, we validate 100 correctly detected malware samples. Results show that 81 out of 100 samples are correctly classified into their respective categories. Table III summarizes the sample distribution and classification accuracy across categories. Due to their less distinct malicious patterns, Adware and Riskware pose greater challenges to accurate analysis compared to other categories, contributing to their higher misclassification rates.
TABLE III: The quality evaluation of generated detection analysis. The “Correct” column indicates the number of categories correctly classified by the LLM.
@@
PUA umpay, scamapp, apptrack 13 9
@@
4.4.3 Cost
We estimate the overall cost of evaluating LAMD. Using GPT-4o-mini as our base LLM, which costs \$0.15 per 1M tokens for input and \$0.6 per 1M tokens for output [46], the total expense for our test set of 9,046 APKs amounts to approximately \$1800, averaging \$0.199 per apk. The cost for each apk varies depending on the number of sliced CFGs extracted and suspicious APIs the app invokes. While this demonstrates the feasibility of deploying LAMD in real-world scenarios, the operational expense remains non-trivial for large-scale applications. Future work could explore the effectiveness of applying LAMD with locally deployed LLMs like DeepSeek [56] and Llama [57] to avoid cloud-based API fees.
@@ 5 Case Study
This section demonstrates how LAMD overcomes the limitations of current LLMs in Android malware detection. Since LLMs cannot process APKs directly, we use JADX [^4] https://github.com/skylot/jadx to decompile them, concatenating all pseudo source codes for input [52]. Besides GPT-4o-mini, we select Gemini 1.5 pro [6] as another comparison, which claims their longest context windows and for its malware detection capabilities.
@@ 6 Discussion
While LLMs improve Android malware detection, their general pre-training limits fine-grained behavior analysis. The phenomenon of over-squashing [40] in LLMs still introduces challenges to successfully identify key aspects within complex programs. Future research can benefit from domain-specific fine-tuning or external knowledge integration [58]. Meanwhile, learning-based methods remain essential for capturing overall insights in non-drift scenarios. Our experiments find that while LLMs excel in adapting to unseen threats, they require significantly more computational resources and incur higher costs, highlighting effective strategies to combine these two paradigms and maximize their complementary strengths. In future work, we plan to explore how LAMD and more generally context-aware LLMs perform against recent approaches aimed at learning temporally invariant features [59, 60]. These findings open interesting research opportunities to better understand the tension between learning-based models and pre-trained LLMs.
@@ 7 Conclusion
Appendix A Baseline Models
Drebin [11] detects malware using binary feature vectors derived from nine data types (e.g., hardware components, API calls, permissions) and classifies samples via a linear classifier.
DeepDrebin [12] extends Drebin by replacing the linear classifier with a three-layer deep neural network (DNN) while retaining the same feature space for feature extraction and classification.
Malscan [45] employs a graph-based approach, extracting sensitive API calls from APKs and computing four centrality measures (degree, Katz, proximity, and harmonic wave centralities) as features. We use the optimal feature with a Random Forest classifier, which leads to the lowest overhead.
All of them can be considered as the current state-of-the-art in Android malware detection.
TABLE IV: Overview of the evaluation dataset, where M denotes malware and B denotes benign applications.
Test Set Time Interval Sample Size Existing Families New Families Packed Malicious (M) Benign (B) M/(M+B) %
Test Set 1 2020.05 – 2021.01 3015 21 24 18 284 2731 9.42
Test Set 2 2021.01 – 2021.12 3015 28 32 30 298 2717 9.88
Test Set 3 2021.12 – 2023.12 3016 34 36 40 302 2714 10.01
TABLE V: Overview of the OOD evaluation dataset aligned with the GPT‑4o‑mini pre-training cutoff (October 2023), where M denotes malware and B denotes benign applications.
Test Set Time Interval Sample Size Existing Families New Families Packed Malicious (M) Benign (B) M/(M+B) %
Learning-based 2020.05 – 2020.06 884 15 7 2 62 822 7.01
LAMD 2023.11 – 2023.12 60 0 1 0 7 53 11.67
TABLE VI: Performance of learning-based methods using Drebin features under different training/validation data volumes.
Sample Size Model Validation Test 1 Test 2 Test 3
@@ Appendix A Baseline Models
Appendix B Impact of Training Dataset Sizes
@@
Intuitively, increased training samples should enhance detection performance in learning-based methods. While LAMD excels in zero-shot learning, one may argue that traditional methods are hindered by smaller training datasets compared to the vast pretraining data leveraged by LLMs. However, the impact of significant drift warrants reconsideration. Table VI highlights the performance of learning-based detectors (e.g., Drebin feature space) across varying dataset sizes, across the training dataset timeframe. The results indicate that additional training data does not improve detection of samples with significant drift, though it enhances performance on in-distribution samples.
@@
Appendix C Factual Consistency Verification
To check factual consistency, we leverage data dependencies that capture relationships between variables and APIs. We focus on five dependencies (see Table VII) as they represent fundamental program relationships critical for malware reasoning and detection. Specifically:
@@ Appendix C Factual Consistency Verification
TABLE VII: Data dependencies used for factual consistency verification.
Category Type Description
Transitive A variable is propagated through a call chain to an API.
Derived A variable is derived from another through computation.
@@
Appendix D Backward Slicing Algorithm